OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 3x – 4 A. (11, 5) and (2, 2) B. (5, 11) and (3, 2) C. (5, 11) and (2, 3) D. (5, 11) and (2, 2) @surjithayer @cutiemarklove

OpenStudy (anonymous):

plug the value of x and find the value of y as I have done previously

OpenStudy (anonymous):

i give you some hint. when x=11 y=3*11-4=33-4=27 so pair is (11,27) but it is not (11,5) so A is rejected. now similarly proceed for B if one pair is true then check for other pair also. if first pair is false then reject it as i have rejected.